Define a filter in time series analysis. Find the effect of a linear filter on the time series

Xt=cosθt+isinθt



1
Expert's answer
2021-12-29T16:34:47-0500

Filtering in time series mean removal of spectral power at certain chosen frequencies while retaining others.

A high-pass filter retains higher frequencies while removing low frequencies; a low-pass filter does the opposite. A band-pass filter removes all frequencies outside a prespecified band.

for Xt =cos"\theta t+ sin\theta t"

filtering involves convolution between two series

X[t]=cos"\theta [t] + sin\theta [t]"

the convolution of the series is the sum of the product between series forward and the series going backwards

X[t] = {cos"\theta" + sin"\theta" }t "\forall t"

introduce the product in both left hand side and right hand side .

"\\\prod X" [t]="\prod" {cos"\theta" +sin"\theta" }t "\forall t"